What is the range of the function f(x)= x^2-4x-12
Answer by lwsshak3(11628)

What is the range of the function f(x)= x^2-4x-12
**
To find the range, interchange x and y, then solve for y.
x=y^2-4y-12
y^2-4y=x+12
completing the square
(y^2-4y+4)=x+12+4
(y-2)^2=x+16
Take sqrt of both sides
y-2=±√(x+16)
y=±√(x+16)+2
radican≥0
x+16≥0
x≥-16
Range of f(x):
[-16,∞)